OpenStudy (anonymous):

11-2x=-7+11 use variable terms and constant terms on both sides of the equation,use additional principal to get variable terms on one side and all constant terms on one side ,then use multiplication principle

OpenStudy (anonymous):

its still 7/2

OpenStudy (anonymous):

-7/2

OpenStudy (anonymous):

-7/-2 = 7/2

myininaya (myininaya):

11-2x=4 -2x=4-11 -2x=-7 2x=7 x=7/2

OpenStudy (anonymous):

correct answer is 0

myininaya (myininaya):

no it is not

myininaya (myininaya):

not for this problem

OpenStudy (anonymous):

no...its its not its 7/2...

myininaya (myininaya):

did you post the correct problem?

myininaya (myininaya):

maybe you should have 11-2x=-7x+11 add 7x on both sides 11+5x=11 subtract 11 on both sides 5x=0 divide 5 on both sides x=0 but you have 11-2x=-7+11